OpenStudy (anonymous):

would someone please help me with this ? Find the range of the function g(x) =3x squared with domain D ={-1,0,1,2,3}

OpenStudy (anonymous):

(3x)^2 or 3(x^2)?

OpenStudy (anonymous):

I'd like to help, but I need to know what the exact function is first. (3x)^2 = 9(x^2), which is different from 3(x^2)

OpenStudy (anonymous):

so it's  based on that, for each value of the domain, you have this : R = { g(-1), g(0), g(1), g(2), g(3)} I'll do one for now, let's see if you can get the rest : g(-1) = 3*(-1) - (-1)^2 = -3 - 1 = -4